### v1

# MapleGrasp: Mask-guided Feature Pooling for Language-driven Efficient Robotic Grasping

Abstract
[…]
Figure 1: Modular systems generate object masks and grasp poses, utilizing a control algorithm to estimate joint space trajectories, while VLAs directly map vision and textual modalities to actions.
1 Introduction
The ability to perceive and interact with objects in an environment is a defining characteristic of human intelligence. As humans mature, we develop capabilities to manipulate novel objects, generalizing our understanding beyond previous experiences. Replicating this intelligence in robotics remains a formidable challenge. Robots often encounter unfamiliar objects and must adapt to new scenarios without prior knowledge [29]. Foundational models, particularly Vision-Language Models (VLMs), aim to bridge this gap by leveraging extensive pre-trained multi-modal representations to enhance robotic perception and interaction [25, 64].
An end-to-end robotic manipulation framework typically comprises four key stages: task planning, visual grounding, object manipulation, and goal-oriented placement. Recent advances have demonstrated the effectiveness of VLMs across these stages, incorporating large-scale pre-trained representations to improve generalization [54]. An alternative paradigm involves Vision-Language-Action (VLA) models, which directly predict trajectories for grasping and manipulating objects, bypassing the modular pipeline [12, 4]. We highlight the difference in these two frameworks in Fig. 1. While VLAs have shown remarkable performance in long-range tasks across both simulation and real-world settings, their deployment in novel environments presents three key challenges: i) Task-Specific Data Collection: Collecting teleoperated or human demonstrations is often difficult and time-consuming for novel interactions [53]; ii) Expensive fine-tuning: Large-scale VLAs require task-specific fine-tuning, which is computationally expensive. Although parameter-efficient fine-tuning (PEFT) techniques alleviate some costs, they restrict adaptability by modifying only a small subset of parameters [20]; iii) Limited generalization: Fine-tuned VLAs often fail to generalize across dynamic environments, particularly when objects are repositioned or when visually similar distractor objects are introduced [46]. Given these challenges, modular frameworks offer an attractive alternative that balances generalizability with ease of deployment. By decoupling perception, planning, and manipulation, modular architectures allow state-of-the-art models in each component to be independently improved and integrated as research advances. Furthermore, LLM-agentic frameworks can serve as high-level decision-makers, dynamically invoking perception and manipulation modules for task execution in novel environments [15, 41, 42]. Compounded errors within modular frameworks can be mitigated using closed-loop feedback and error planning [6, 16].
In this work, we focus on a critical capability within modular frameworks, language-driven grasp detection. This task involves predicting a stable grasp pose for a specified object based on RGB-D images and natural language referring expressions. We focus on 2D grasp detection, which can be naturally extended to 3D using depth information and camera intrinsic parameters. Our contributions include:
  1. 1.
    A novel two-stage training strategy for LDRG: initially predicting text-referred object masks, then refining grasp predictions via mask-guided feature pooling.
  2. 2.
[…]
  3. 3.
    Introduce RefGraspNet, a large-scale open-source dataset comprising over 200 million grasp annotations, eight times larger than existing datasets, consisting of challenging real-world scenarios for grasping.
  4. 4.
    Extend 2D grasp predictions to 3D via depth maps and inverse kinematics, achieving performance comparable to recent Vision Language Action (VLA) models on the LIBERO benchmark, with superior generalization to diverse tasks while being parameter efficient.
2 Related Works
Grasp Detection. Robotic grasping with parallel grippers has been extensively studied in both the 2D and 3D domains. The 4-DoF grasp representation defines a grasp using four parameters: $(x,y)$ as the center of the grasp, $w$ as the grasp width, and $\theta$ as the grasp axis angle relative to a top-down hand orientation [1, 8, 17]. The 6-DoF representation extends this to 3D, incorporating three degrees of translation and three degrees of rotation required for stable object grasping [31, 10, 43]. Early approaches to object grasping relied on kinematics and geometric methods to identify stable grasp poses from depth maps [3] and 3D object models, using datasets generated through offline labeled training data, human demonstrations, or trial-and-error methods [13, 9]. The use of RGB-D data has shown promising results for precision grasping, where CNN and ResNet-based architectures leverage visual features to predict grasp pose parameters [22, 48, 65, 58]. Additionally, 3D-CNN models have been employed to extract 3D features from point clouds [39, 38, 51]. Recent advancements in transformer architectures and self-supervised learning have bridged the gap between vision and textual modalities, shifting research towards incorporating language as an additional modality for object manipulation [54]. Language-driven grasp detection (LDGD) refers to the task of associating referring text expressions with RGB-D images to predict a suitable robotic grasp pose for object manipulation [49]. Two primary approaches have emerged for LDGD: i) Visual grounding of referring expressions: This approach predicts segmentation masks for the target object [7, 55, 61], followed by the application of pre-trained grasp detection networks to compute grasp poses [2, 27, 32]. ii) End-to-end grasp pose estimation: This method directly predicts grasp poses from input text and RGB-D modalities [56, 63, 52].
Figure 2: Our framework trains on fused vision-language embeddings for object segmentation, then applies mask pooling in grasp prediction heads. MapleGrasp is an network-agnostic MFP block with no extra parameters and can be used with the CROG backbone as shown.
End-to-End Language-Driven Grasping. Our work focuses on utilizing RGB images of a scene along with free-form referring expressions to determine a stable 4-DoF grasp for the target object in cluttered environments. We extend this by incorporating depth maps to transform the 2D grasp into 3D and leveraging a differential inverse kinematics controller to compute the joint-space trajectory for grasp execution. Several studies have explored similar setups for LDGD [45, 46]. Recent work has introduced diffusion models that learn a probabilistic diffusion process to generate grasp poses during inference [49, 35, 36]. Vision-language fusion provides an alternative approach, offering faster convergence and lower inference times [34, 24, 54]. The ETRG framework enhances vision-language feature fusion for grasp detection by integrating vision-language adapters into frozen CLIP vision and text encoders [57]. MaskGrasp employs a tri-head architecture with cross-attention mechanisms using text embeddings, region-of-interest (ROI) feature embeddings, and mask embeddings [47]. GraspSAM adapts the Segment Anything Model [21] for grasp detection and utilizes GroundingDINO [28] to transform text expressions into object bounding boxes [37].
In our work, we analyze the impact of mask pooling on grasp detection and demonstrate that combining mask pooling with open-set object detectors significantly enhances performance on unseen objects—an issue that previous studies have not adequately addressed. Moreover, we present the first exhaustive comparison of LDRG approaches against large Vision-Language-Action (VLA) models for language-guided grasping tasks.
Vision-Language-Action Models. Direct modeling of vision and language modalities for joint-space robotic control has gained significant interest in recent research. Advances in vision-language models [59, 18, 30], coupled with the availability of large-scale robotic datasets spanning diverse environments, hardware platforms, and tasks [5], have enabled the training of Vision-Language-Action (VLA) models for robotic manipulation [40, 14, 26]. However, the large model sizes of VLAs often demand substantial GPU resources for fine-tuning and real-world deployment [50]. Recent research has sought to mitigate these computational challenges through parameter-efficient training techniques and action chunking methods [19, 20], enabling deployment on smaller GPUs with real-time kernel compatibility for seamless robotic control.
3 Mask-Guided Feature Pooling for Language Driven Grasp Detection
In this section, we introduce our approach of leveraging segmentation masks for feature pooling to focus on relevant regions in an image for targeted grasp pose estimation in cluttered environments. Furthermore, we present our two-stage training framework and discuss our improved training objective used to optimize our method.
3.1 Language-Driven Grasp Detection
Given an RGB-D image and a corresponding referring expression, language-driven grasping models predict a stable grasp pose to pick up the specified object using either a parallel gripper or a dexterous robotic hand. This task poses three primary challenges: i) accurately identifying the target object using textual cues provided by the referring expression, such as color, shape, and relative positioning; ii) determining a stable and collision-free grasp pose that can be reliably executed in real-world environments without interference from surrounding objects; and iii) generalizing grasp predictions effectively to previously unseen objects that were not included in the training dataset.
We focus on grasp pose estimation in 2D, where where grasps are typically represented as rectangles [49, 37, 34]. Each 2D grasp is defined by the center coordinates $(x,y)$ of the grasp rectangle, the grasp width $w$, and the orientation angle $\theta$ relative to a fixed horizontal axis. Given that this study focuses on top-down grasps performed with a parallel gripper, the height of the grasp rectangle remains constant. Such 2D grasp representations naturally extend into 3D space by incorporating depth information.
Formally, a 2D grasp $\mathbf{g_{2D}}=(x,\,y,\,w,\,\theta)$ is converted into a 3D grasp pose $\mathbf{T_{3D}}$ by projecting the pixel $(x,y)$ into 3D space. With depth value $D(x,y)$ at pixel $(x,y)$, the 3D coordinates relative to the camera are computed as:
$Z=D(x,y),\quad X=\frac{(x-c_{x})\,Z}{f_{x}},\quad Y=\frac{(y-c_{y})\,Z}{f_{y}},$ (1)
where $f_{x},f_{y}$ denote the camera focal lengths, and $c_{x},c_{y}$ represent the principal point offsets. The computed coordinates $(X,Y,Z)$ correspond to the grasp center position in the camera frame. The resulting top-down 3D grasp pose transformation matrix $\mathbf{T}_{3D}$ is then defined as:
$\mathbf{T}_{3D}=\begin{bmatrix}\cos\theta&-\sin\theta&0&X\\ \sin\theta&\cos\theta&0&Y\\ 0&0&1&Z\\ 0&0&0&1\end{bmatrix}.$ (2)
Direct prediction of all four grasp parameters (location, width, and orientation) at every pixel often yields suboptimal results due to their weak direct correlation with the original RGB-D data. Thus, contemporary approaches typically estimate grasp parameters using three distinct 2D prediction maps: i) Grasp Quality Map: Assigns a pixel-wise score for grasp quality; the pixel with the highest score represents the best grasp location, ii) Grasp Width Map: Encodes the optimal grasp width at each pixel and iii) Grasp Angle Map: Stores the grasp orientation $\theta$ at each pixel. This decoupled representation enhances the reliability and interpretability of grasp prediction models.
3.2 Referring Expression Segmentation
Language-Driven Grasp Detection (LDGD) shares strong similarities with a well-established computer vision task known as Referring Expression Segmentation (RES). The goal of RES is to predict a binary segmentation mask that separates the referred object from its surroundings [33, 62]. Since both tasks involve object-level visual grounding, several LDGD frameworks incorporate both segmentation mask prediction and grasp rectangle estimation from the referring query [47, 57]. These architectures typically contain separate projection layers for predicting the binary object mask and the grasp maps. However, since predictions are made independently, the segmentation mask may be correctly identified while grasp maps remain inaccurate.
3.3 Motivation for Mask Pooling
The predicted binary segmentation masks highlight the image regions corresponding to the referring expression, providing an implicit constraint for grasp pose estimation. If the segmentation mask is correctly predicted, then the optimal grasp pose should lie within the segmented region. This suggests that the grasp prediction module does not need to consider grasp locations outside the segmented object, provided that the segmentation branch can accurately localize the object. With recent advancements in VLMs for RES, zero-shot object segmentation has significantly improved. Consequently, off-the-shelf segmentation models can be used to extract masks for grasp pose estimation networks to predict stable grasp poses [7, 61]. We hypothesize that incorporating segmentation masks as a guiding mechanism within the grasp detection pipeline can improve accuracy by focusing predictions on relevant regions. Prior work has not explored this direction, and we argue that integrating mask pooling with LDRG can effectively utilize advances in RES to enhance grasp detection accuracy.
3.4 Mask-Guided Feature Pooling (MFP)
The concept of utilizing segmentation masks for pixel pooling can be integrated into various LDGD frameworks that jointly predict segmentation masks and grasp maps. We illustrate this integration using the CROG framework [46], a robust LDGD model that leverages CLIP-based visual and textual encoders combined through cross-attention mechanisms to fuse vision-language features. CROG employs a multi-modal Feature Pyramid Network (FPN) followed by additional cross-attention layers that incorporate textual embeddings. These enriched features are subsequently passed through a multi-task projector, which simultaneously outputs binary segmentation masks and grasp parameter maps. The projector architecture consists of sequential Conv2D and upsampling layers designed to maintain feature dimensions consistent with the original input images.
We introduce Mask-Guided Feature Pooling (MFP) within the multi-task projector (see Fig. 2). Our training methodology comprises two sequential stages: i) Pretraining the Segmentation Mask Branch: In this initial stage, we freeze the grasp prediction heads and train the segmentation mask branch separately under a standard Referring Expression Segmentation (RES) training setup. This stage intentionally allows slight overfitting to ensure high-quality object-level masks, facilitating precise pixel pooling in subsequent training. ii) Mask-Guided Grasp Prediction: In this stage, we introduce mask-guided pooling connections within the multi-task projector, focusing the grasp predictions exclusively on visually grounded object regions defined by the segmentation masks. These pooled features undergo further refinement via a simple Multi-Layer Perceptron (MLP) to generate accurate grasp maps.
Two critical design choices ensure robustness. First, we do not use ground-truth object masks for pooling, as this would make grasp predictions overtly sensitive to mask accuracy, which is challenging in cluttered scenes with complex referring text. Second, we avoid using open-set detectors during training, since state-of-the-art object mask segmentation models currently achieve only 50–60% IoU on our benchmarks, which is insufficient for reliable pooling.
3.5 Loss Function
Prior work utilizes the Smooth L1 loss for pixel-wise prediction in visual grounding, penalizing smaller errors quadratically and larger errors linearly, defined per pixel $(\hat{x}_{p},x_{p})$ as:
$L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})=\begin{cases}\tfrac{(\hat{x}_{p}-x_{p})^{2}}{2\beta},&\!\text{if }|\hat{x}_{p}-x_{p}|<\beta,\\ |\hat{x}_{p}-x_{p}|-\tfrac{\beta}{2},&\!\text{otherwise.}\end{cases}$ (3)
We extend this to formulate weighted Smooth L1 loss $w_{p}\,L_{\mathrm{smoothL1}}(\hat{x}_{p},x_{p})$, by assigning per‐pixel weights $w_{p}=1+\alpha\,Q_{\mathrm{gt},p}$, where $Q_{\mathrm{gt},p}\in[0,1]$ is the ground‐truth quality map indicating the object grasp region. By emphasizing suitable grasp regions, we enhance precise grasp estimation. Our experiments demonstrate that this modified loss function is beneficial in low-resource settings, where the dataset contains only single grasp annotation per object, ensuring more reliable and efficient learning.
3.6 Inverse Kinematics Control for Manipulation
Our approach for real world object manipulation transforms the 2D grasp prediction into a desired 3D pose. Specifically, we obtain the final grasp pose in the robot’s base frame, $\mathbf{T}_{\text{base}}^{\text{pose}}$, by composing three transformations:
$\mathbf{T}_{\text{base}}^{\text{pose}}\;=\;\mathbf{T}_{\text{base}}^{\text{ee}}\;\mathbf{T}_{\text{ee}}^{\text{cam}}\;\mathbf{T}_{\text{cam}}^{\text{pose}},$ (4)
where:
  1. 1.
    $\mathbf{T}_{\text{base}}^{\text{ee}}$ is the forward-kinematics transformation from the robot base to its end-effector,
  2. 2.
    $\mathbf{T}_{\text{ee}}^{\text{cam}}$ is the hand–eye calibration matrix (mapping from end-effector to camera),
  3. 3.
    $\mathbf{T}_{\text{cam}}^{\text{pose}}$ is the camera-to-pose transformation predicted by our grasp model.
To execute a grasp in closed loop, we generate a smooth interpolation $\mathcal{C}(t)$ in $\mathrm{SE}(3)$ between the initial end-effector pose $\mathbf{T}_{\text{init}}$ and the desired grasp pose $\mathbf{T}_{\text{goal}}$, as:
$\mathcal{C}(t)\;:\;[0,\,1]\;\to\;\mathrm{SE}(3),\quad\mathcal{C}(0)=\mathbf{T}_{\text{init}},\quad\mathcal{C}(1)=\mathbf{T}_{\text{goal}}.$ (5)
At each control step $t$, the desired pose is
$\mathbf{T}_{\text{des}}\;=\;\mathcal{C}(t).$ (6)
Let $\mathbf{T}_{\text{current}}$ denote the current pose of the end effector. We define the positional error
$\mathbf{e}_{p}=\mathbf{p}_{\text{des}}-\mathbf{p}_{\text{current}},$ (7)
and compute the rotational error using the matrix logarithm of the orientation difference:
$\mathbf{R}_{\text{err}}=\mathbf{R}_{\text{des}}\,\mathbf{R}_{\text{current}}^{\top},\qquad\boldsymbol{\theta}_{\text{err}}=\mathrm{Log}\!\bigl{(}\mathbf{R}_{\text{err}}\bigr{)},$ (8)
where $\mathbf{R}_{\text{des}}$ and $\mathbf{R}_{\text{current}}$ are the rotation components of $\mathbf{T}_{\text{des}}$ and $\mathbf{T}_{\text{current}}$, respectively, and $\mathrm{Log}(\cdot)$ is the matrix logarithm on $\mathrm{SO}(3)$. We form the combined error vector
$\mathbf{e}=\begin{bmatrix}\mathbf{e}_{p}\\ \boldsymbol{\theta}_{\text{err}}\end{bmatrix}\;\in\;\mathbb{R}^{6}.$ (9)
The desired end-effector velocity is then
$\dot{\mathbf{x}}_{\text{target}}\;=\;K_{p}\;\mathbf{e},$ (10)
where $K_{p}$ is a diagonal gain matrix. We compute the joint velocity command by projecting $\dot{\mathbf{x}}_{\text{target}}$ through the robot Jacobian pseudoinverse:
$\dot{\mathbf{q}}_{\text{cmd}}\;=\;\mathbf{J}^{\dagger}\;\dot{\mathbf{x}}_{\text{target}}.$ (11)
Finally, we update the robot’s joint states with $\dot{\mathbf{q}}_{\text{cmd}}$ and advance $t\!\leftarrow\!t+\Delta t$. This process repeats until
$\|\mathbf{T}_{\text{des}}-\mathbf{T}_{\text{current}}\|\;\leq\;\epsilon,$ (12)
indicating successful convergence to the target grasp pose.
4 RefGraspNet Dataset: Learning to Grasp from 200M+ Grasp Poses
Dataset Real/Simulated No. of Images Ref. Text Unique Obj. Grasp Poses Obj. Masks Avg. Obj./Img.
Jacquard [8] 54K 967K 1
VMRD [60] 4.5K 31 51.5K 3.5
OV-Grasp [23] Both 63K 117
ACRONYM [10] 262 17.7M
RoboRefIt [32] 10.7K 50.7K 66 7.1
OCID-VLG [46] 1.7K 89.6K 31 521K 17.08
GraspAnything++ [49] 994K 10M 236 33M 3.4
Ours: RefGraspNet 97K 12.25M 88 219M 9.57
Table 1: Comparison of existing 4-DoF LDRG datasets. RefGraspNet provides 200M+ high quality grasps, object masks for each scene, and thus can be used to finetune both referring segmentation and grasping models.
We build on existing research and datasets in language driven grasping by creating a new corpus of real world cluttered images and their grasp annotations. Our dataset, RefGraspNet, is derived from the GraspNet-1B dataset [11] and comprises a diverse corpus collected using 88 unique objects arranged in cluttered scenes. We construct 10 instruction templates such as “Grasp the {object}”, “Find the {object} in this scene”, etc. and expand the object names to create referring expressions. In scenes with multiple objects, we leverage an open-source model, DeepSeek-VL [30], to provide additional attributes such as color, shape and relative position. We chose GraspNet-1B as our base corpus since it consists of 1 billion + grasp poses annotated with sampling and force-closure techniques. We sample grasps with a confidence score greater than 70%, and our final corpus comprises 219M grasp poses, 8 times bigger than the previous largest dataset for language driven 4 DoF grasping. Comparisons with prior datasets are provided in Tab. 1. We randomly choose 70% of the objects to be seen objects (with the rest as unseen) and create train/val/test-seen and test-unseen splits for our experiments.
5 Experiments
Here, we present our experiments evaluating MapleGrasp on OCID-VLG and RefGraspNet benchmarks, while also comparing our framework to prior LDRG and VLA methods in simulation and real-world setups.
5.1 OCID-VLG Benchmark
The OCID-VLG corpus [46] contains 1763 highly cluttered indoor tabletop RGB-D scenes with 31 unique objects. Objects have annotated 4-DoF grasp poses and segmentation masks, associated with referring expressions describing attributes such as color, shape, and relative position. With 89K unique (RGB-Text-Mask) tuples, it is the largest publicly available benchmark for language-driven grasp detection frameworks [^1] GraspAnything++ is another open-source corpus, but the publicly shared subset is significantly smaller than what was used for training their models, preventing fair benchmarking.. We evaluate various CLIP-based architectures on OCID-VLG, applying the MapleGrasp framework to Ref2Grab [2], which uses frozen CLIP encoders for improved generalization, and CROG [46], as detailed in Sec. 3.4. Results are summarized in Tab. 2.
Model IOU Pr@50 Pr@60 Pr@70 Pr@80 Pr90 J@1 J@Any
DetSeg + CLIP 29 27.2 20.9 17.5 17.2 16 28.1 39.2
GR-ConvNet + CLIP 31.3 21.0 11.6 5.5 2.4 0.5 9.73 15.41
SSG + CLIP 33.6 35.6 35.6 35.5 35.5 32.8 33.5 34.7
CROG 81.1 96.9 94.8 87.2 64.1 16.4 77.2 87.7
Ref2Grab 83.26 85.68 85.13 91.53 89.69 83.21 70.5 79.1
ETRG 80.11 - - - - - 82.28 91.12
MapleGrasp-Ref2Grab 83.78 85.96 85.41 91.74 90.13 83.79 76.8 84.7
MapleGrasp-CROG 81.36 97.40 95.32 87.9 65.4 16.4 86.15 91.9
Table 2: Results on OCID-VLG. We implement all baseline methods except ETRG, whose results are taken from their paper. Our comparison is limited to methods that do not utilize depth maps during training.
Referring Text Molmo+SAM+GraspNet CROG MapleGrasp
Grasp the rightmost red food box [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Grab the transparent food bag [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Find the banana furthest away from me [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 3: Qualitative comparisons of LDRG approaches under complex referring queries. The zero-shot baseline (first column) struggles to identify referred object due to distractors. MapleGrasp (last column) can predict a stable and accessible grasp for each queried object.
Applying mask proposals enhances both the CROG and Ref2Grab baselines in terms of the Jacquard Index J@N metric, which evaluates the top-N grasp rectangle candidates that exhibit a rotation angle difference of less than $30^{\circ}$ and an IoU greater than 0.25 compared to the ground-truth grasp rectangle. MapleGrasp achieves a 12% relative improvement in J@1 over previous state-of-the-art methods, demonstrating higher precision in grasp prediction. However, referring expression segmentation (RES) accuracy, measured using Prec@X, which quantifies the percentage of test cases where the IoU exceeds $X\in\{0.5,0.6,0.7,0.8,0.9\}$, remains relatively unchanged. This may be due to CROG losing pretrained knowledge during CLIP layer fine-tuning. In contrast, MapleGrasp-Ref2Grab, which employs frozen vision-language encoders, achieves higher performance in the RES task but struggles with grasp map prediction due to its limited trainable parameter space (4M vs. CROG’s 167M). Overall, our findings suggest that MFP enhances both frozen and fully fine-tuned CLIP-based models for language-driven grasp detection.
5.2 Ablation Studies
We conduct ablation studies on MapleGrasp (LABEL:tab:ablations) to analyze its effects due to architectural modifications. Results highlight the importance of accurate RES for mask pooling; using Molmo, which achieves lower IoU on RES tasks, negatively impacts grasp accuracy due to complex referring queries requiring task-specific visual grounding. While using frozen CLIP improves RES, it decreases grasp performance because of fewer trainable parameters. Replacing cross-attention with MLP-Mixers, despite reported benefits [44], does not significantly improve performance. Furthermore, replacing the weighted smooth L1 loss with a standard Smooth L1 loss leads to reduced scores, likely from slower convergence. All ablations use a fixed number of training epochs for consistency.
Method IoU J@1 J@Any
MapleGrasp-CROG 81.36 86.15 91.90
(1) Molmo+SAM for Mask Pooling 73.42 77.79 80.16
(2) Frozen CLIP Layers 83.52 72.19 74.44
(3) No Cross-Attn, MLP-Mixer 80.16 84.37 90.45
(4) Standard Smooth L1 Loss 80.17 83.23 89.18
5.3 RefGraspNet Benchmark
We further evaluate previous baselines and MFP-enhanced architectures on our RefGraspNet dataset, establishing a strong benchmark for future research. The dataset comprises 88 unique objects, with 70% assigned as seen objects and 30% as unseen objects. Results for both test splits are reported in Tab. 4. Similar trends as observed in the OCID-VLG benchmark hold here, with MFP providing significant performance gains for both Ref2Grab and CROG. However, all trained models perform worse on unseen objects, highlighting a lack of zero-shot generalization. In contrast, open-set object detectors such as Molmo [7] achieve higher test unseen IoU. This motivates us to incorporate Molmo in an ensembling setup to boost zero-shot segmentation, which is essential for robust grasp pose estimation.
Test Seen Test Unseen
Model IoU J@1 J@Any IoU J@1 J@Any
DetSeg + CLIP 48.18 40.19 41.56 30.57 27.18 27.20
GR-ConvNet +
CLIP 52.34 44.19 46.56 32.18 30.19 32.17
SSG + CLIP 57.16 48.78 50.10 38.19 34.16 34.89
CROG 84.49 85.32 86.49 69.95 70.81 71.99
Ref2Grab 87.32 73.27 75.45 75.67 62.12 63.28
MapleGrasp-Ref2Grab 87.74 75.67 77.89 75.11 65.19 67.22
MapleGrasp-CROG 84.86 89.15 89.67 70.43 74.92 75.67
Table 4: Results on RefGraspNet. The MFP framework leads to consistent improves across seen and unseen objects. However, performance on unseen objects deteriorates for all methods.
5.4 Cross-Dataset Testing
To assess the generalizability of models trained on RefGraspNet, we cross-test them on the OCID-VLG benchmark. We compare MapleGrasp-Ref2Grab and MapleGrasp-CROG, and present results in Tab. 5. Our findings reveal that models trained on RefGraspNet generalize better than those trained solely on OCID-VLG. The dataset’s diversity in grasp poses and text expressions enhances inference on other benchmarks. This highlights the advantages of training on RefGraspNet, encouraging further adoption of our corpus to push performance boundaries.
OCID-VLG RefGraspNet
Model Train-Dataset J@1 J@Any J@1 J@Any
MapleGrasp-Ref2Gr OCID-VLG 76.8 84.7 41.8 42.9
RefGraspNet 71.2 74.4 75.7 77.9
MapleGrasp-CROG OCID-VLG 86.2 91.9 33.2 35.6
RefGraspNet 68.2 73.8 89.2 89.7
Table 5: Cross-dataset evaluation results. Models trained on RefGraspNet generalize better than OCID-VLG.
5.5 LIBERO Simulation Experiments
We deploy the MapleGrasp model trained on RefGraspNet and extend grasp poses to 3D using an inverse kinematics controller for trajectory prediction (Sec. 3.6). This setup allows direct comparisons with Vision-Language-Action (VLA) models. Since our approach focuses on language-driven grasp pose estimation, we cannot evaluate task success rates directly. Instead, we define grasp success as trials in which an object is lifted and held above ground level for at least two seconds. We report J@1 and J@Any metrics. Experiments are conducted on three task suites from the LIBERO benchmark: SPATIAL, GOAL, and OBJECT. Since VLAs are fine-tuned on 50 teleoperated demonstrations per task, we automate 4-DoF grasp annotations using forward kinematics-based 3D grasp pose estimation and project the masks to 2D, inverting steps described in Sec. 3.1. We fine-tuned MapleGrasp on these annotations to adapt to the simulation environment. At inference time, MapleGrasp predicts 2D grasps, which are transformed and executed in 3D by the controller. Results in Tab. 6 compare our method with VLA and diffusion baselines.
Model SPATIAL GOAL OBJECT
Diffusion Policy 0.77 0.69 0.92
Octo 0.78 0.84 0.88
OpenVLA 0.84 0.79 0.90
MapleGrasp-CROG 0.87 0.85 0.90
Table 6: Libero simulation results across SPATIAL, GOAL and OBJECT task suites. Scores are reported using J@1.
Our results demonstrate that our grasping+control pipeline achieves similar accuracy to recent VLAs in grasping tasks. Interestingly, decoupling grasping and control improves generalization, as trajectory-based imitation learning methods fail to complete unseen tasks. However, our model successfully determines grasp poses for similar but unseen objects, allowing the control algorithm to generate a stable trajectory without requiring video demonstrations. Cross-dataset testing in Tab. 7 reveals the capabilities of our approach in handling unseen queries across task suites. Since our LDRG model is only concerned with predicting grasp poses and we leverage a controller to determine the trajectory, our method can be deployed without learning from tele-operated demonstrations.
Model Train-Dataset GOAL OBJECT
OpenVLA GOAL 0.79 0.00
OBJECT 0.00 0.90
MapleGrasp-CROG GOAL 0.85 0.62
OBJECT 0.68 0.90
Table 7: Comparison of OpenVLA and our method in cross-dataset settings. OpenVLA struggles with seen objects under novel tasks, while our method generalizes well to unseen queries.
5.6 Real-Robot Experiments
Figure 3: Setup for robot trials using 7 DoF franka arm and our modular pipeline described in section Sec. 3.6
We evaluate various LDRG methods in a real-world robotic setup using a Franka arm with a mounted camera at the gripper, as shown in Fig. 3. We construct diverse tabletop scenes featuring single and multiple objects, distractors, and varying lighting conditions. Results are presented in Tab. 8. For real-robot trials, MapleGrasp integrates Molmo-based object mask refinement, which compares our model’s segmentation predictions with Molmo’s out-of-vocabulary predictions, selecting the highest mutual overlap. MapleGrasp outperforms other baselines, achieving a notable 7% relative improvement in cluttered setups.
Model Single Cluttered
Det-Seg+CLIP 0.27 0.21
GR-ConvNet+CLIP 0.30 0.27
Molmo+SAM+GraspNet 0.45 0.47
SSG+CLIP 0.21 0.17
Ref2Grab 0.42 0.40
CROG 0.55 0.53
MapleGrasp-CROG 0.59 0.57
Table 8: Real robot experiments. Success rate measured across 10 trials for every grasp instruction.
6 Conclusion
Our results highlight mask-guided feature pooling as a simple yet effective technique for language-driven robotic grasping. By selectively pooling predictions within target regions, MapleGrasp reduces misidentifications and grasp errors in cluttered environments. Our two-stage training strategy, which first predicts segmentation masks and subsequently refines grasp predictions, further enhances accuracy and computational efficiency. Compared to end-to-end vision-language-action models, our modular approach demonstrates superior scalability and robustness, and enables rapid adaptation using readily available vision-language models. Moreover, these systems eliminate the need for time-intensive teleoperated demonstrations and can be customized using a small corpus of RGB-text-grasp annotations. Real-world validation confirms significant improvements in grasp accuracy and efficiency.